# Taxonomy of differentiationP. Komarov, F. van Breugel, J. N. Kutz

### v1

@@ 7.4.1 Sliding Window Polynomial Fits
A function $x(t_{i})=c_{0}+c_{1}t_{i}+...c_{d}t_{i}^{d}$, where $d$ is some degree, can be fit to data $\mathbf{y}$ by solving the linear inverse problem:
@@ 7.4.2 Savitzky-Golay Filtering
This is seen by an argument very similar to the derivation of Finite Difference equations in section 4.1. Specifically, the moving window implies a stencil of $S$ samples, $s_{0},s_{1},...s_{S-1}]$, e.g. [1,0,1][-1,0,1] to sample 3 equispaced points centered on a point of interest. We then wish to find a $d^{\text{th}}$ order polynomial of the form $c_{0}+c_{1}s+c_{2}s^{2}+...c_{d}s^{d}$ that when sampled at the stencil-points approximates the data, $\mathbf{y}_{\text{window}}$, at the corresponding points. Paralleling Equations 2 and 39, this can be set up as:
@@
$\frac{d}{dt}(c_{0}+c_{1}s+...c_{d}s^{d})=\frac{d}{ds}(c_{0}+c_{1}s+...c_{d}s^{d})\frac{ds}{dt}=\frac{c_{1}}{\Delta t}$
@@ 7.4.3 Spline Smoothing
where $S$ is the overall spline, $\boldsymbol{\theta}$ represents all coefficients collected together, and $\lambda$ is a hyper-parameter which balances fitting the data with keeping the second derivative small. The second possibility [dierckx1975] is to simply require the sum of squares of the best fit be within a bound, $s$, which corresponds to the minimization of Equation 40 with $\lambda=0$ and an additional condition:
@@ 7.5.2 Radial Basis Functions
Local functions with independent variable given by a radial distance, $r$, which evaluate to zero beyond some maximum distance, $\rho$—such as truncated Gaussians, $e^{-r^{2}/(2\sigma^{2})},\ r<\rho$—can also constitute a linearly independent basis, making the matrix in the linear inverse problem of Equation 5 banded, symmetric, and invertible in $O(N\rho^{2})$. We call this the $\mathbf{A}$ matrix, and visualize an example in the first panel of Figure 42. Unfortunately, $\mathbf{A}$ is typically ill-conditioned, having eigenvalues very close to 0, which makes coefficient solutions highly sensitive to noise in the target, $\mathbf{y}$. Unlike the Fourier case where sensitive higher modes corresponding to tiny eigenvalues of $\mathbf{A}$ can be zeroed out, discarding any element of a local basis significantly deforms function reconstruction in some region, so we need to keep them all.
@@
By taking the gradient w.r.t. $\mathbf{c}$ and setting equal to 0, the above is minimized by solving the linear inverse problem $(\mathbf{A}^{T}\mathbf{A}+\lambda\mathbb{I})\mathbf{c}=\mathbf{A}^{T}\mathbf{y}$. We can eigendecompose symmetric $\mathbf{A}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{T}$ with $\mathbf{V}^{T}\mathbf{V}=\mathbb{I}$, and manipulate the solution into $(\boldsymbol{\Lambda}+\boldsymbol{\Lambda}^{-1}\lambda)\mathbf{c}^{\prime}=\mathbf{y}^{\prime}$, where $\mathbf{c}^{\prime}=\mathbf{V}^{T}\mathbf{c}$ and $\mathbf{y}^{\prime}=\mathbf{V}^{T}\mathbf{y}$ are just vector rotations. Because the matrices are diagonal, we can find $c_{i}^{\prime}=y_{i}^{\prime}/(\lambda_{i}+\lambda/\lambda_{i})$. In words, this means that for large eigenvalues, rotated coefficients are approximately proportional to rotated data divided by those eigenvalues, and as eigenvalues tend toward zero, coefficients do likewise.
@@
To see what this is doing relative to Tikhonov, we again eigendecompose to find $(\boldsymbol{\Lambda}+\mathbb{I}\lambda)\mathbf{c}^{\prime}=\mathbf{y}^{\prime}$, which has component-wise solution $c_{i}^{\prime}=y_{i}^{\prime}/(\lambda_{i}+\lambda)$. In words, large eigenvalues cause coefficients to be inversely proportional as before, but as $\lambda_{i}\to 0$, $c_{i}^{\prime}\to\frac{1}{\lambda}$ rather than 0 for Tikhonov or $\infty$ in the undamped case.
@@
$\hat{x}(t)=\sum_{k=0}^{N-1}c_{k}\xi_{k}(t),\quad\xi_{k}(t)=\begin{cases}e^{-\frac{(k-t)^{2}}{2\sigma^{2}}}&|k-t|<\rho\\ 0&\text{otherwise}\end{cases}$
@@ 7.8.2 Accuracy and Bias
The other major consideration is method accuracy in terms of RMSE (Equation 33) and bias in terms of Error Correlation (Equation 34) against the true, withheld derivative. We are interested in whether the choice of method should change with different data characteristics: outliers, noise type, noise level, and the size of (consistent) $\Delta t$. We are also interested in whether the idiosyncrasies of particular data series favor particular differentiation algorithms, and we would like to know how sensitive methods are to the choice of cutoff frequency (bandlimit), which affects the smoothness of the optimized answers through the loss function’s $\gamma$, set according to Equation 37 heuristic. In the presence of outliers, we optimize the robust loss function in Equation 36 with Huber parameter $M=2$, and in the absence of outliers choose M=6M=6.272727We have observed empirically that, under conditions of Gaussian noise, optimizing the robust loss with M=6M\!=\!6 essentially always produces identical answers to optimizing the original RMSE-based loss, Equation 35. This makes sense theoretically, because inlier errors beyond $6\sigma$ from the mean error are exceedingly rare: 2*(1-scipy.stats.norm.cdf(6)) = 1.97e-9. Lowering $M$ to 2 provides a slight performance edge in the presence of outliers, allowing them to be weighted lighter, but in the absence of outliers M=2M\!=\!2 does begin to degrade performance—yet still only very slightly, because the portion of inliers within $2\sigma$ is more than 95%.
@@
The analysis space now has 7 dimensions: differentiation method, simulation, step size, noise type, noise scale, whether there are outliers, and cutoff frequency. There is also a bonus dimension: To get a sense of performance center and scatter, we need several rounds of simulation followed by optimization with different random seeds. Some dimensions must be explored exhaustively: 12 methods, 6 simulations, and we turn the number of random seeds all the up to 52. To save compute and target our analysis at specific questions, we sweep along slices of the remaining orthotope, only varying one parameter at a time. We choose slices to all pass through a common central point, {cutoff frequency=3 Hz, $\Delta t$=0.01, noise type=normal, noise scale=1, outliers=False}), as symbolized in Figure 51.
@@ 7.9 Recommendations
To further compare methods, auxiliary considerations are summarized in Figure 57. These factors emphasize the value of algorithm flexibility, which can be underappreciated in theoretical discussion but is often important in practice. Indeed, one of the reasons neural networks have become so dominant is their endless capacity for recombination into architectures tailored to multivariate, multidimensional, sequential, incomplete, and structured data.292929“Time cannot wither nor custom stale her infinite variety.” –Shakespeare on the perceptron We recommend RTSDiff (section 7.7) as a general-purpose method for its outstanding versatility and narrowly superior accuracy in many experiments.
@@
Figure 57: Ability of methods to handle auxiliary considerations. Black tildes mean neither yes or no, often that a method could be extended to handle a situation, but at the cost of computational efficiency or mathematical or code simplicity. The heavy dependency at issue is a set of convex solvers, which, although easy to install and use in modern form [cvxpy], are based on underlying compilers and iterative algorithms that can fail to solve efficiently or at all, a source of potential brittleness.
@@ 8 Irregularly-Spaced Samples
[…]
@@ 8.1 Finite Difference with Irregular Steps
There is nothing stopping us from solving the stencil Vandermonde inverse problem of Equation 2 with irregular stencils. We do this by multiplying both sides by $\Delta x$ to put stencil locations back in units of the independent variable and then using whichever fractional or decimal stencil locations, $s_{0},...s_{S-1}]$, match data locations.
@@ 8.3 Kalman Smoothing with Irregular Steps
$=\mathbb{E}\big[\big(e^{\mathcal{A}\Delta t}\mathbf{x}(0)+\cancel{\mathbf{B}(\Delta t)\mathbf{u}}+\mathbf{w}_{d}(\Delta t)-e^{\mathcal{A}\Delta t}\mathbb{E}[\mathbf{x}(0)]$
$\quad\quad\quad\quad-\cancel{\mathbf{B}(\Delta t)\mathbf{u}}-\mathbb{E}[\mathbf{w}_{d}(\Delta t)]\big)\big(\text{ditto}\big)^{T}\big]$

### v2

@@ 7.4.1 Sliding Window Polynomial Fits
A function $x(t_{i})=c_{0}+c_{1}t_{i}+...+c_{d}t_{i}^{d}$, where $d$ is some degree, can be fit to data $\mathbf{y}$ by solving the linear inverse problem:
@@ 7.4.2 Savitzky-Golay Filtering
This is seen by an argument very similar to the derivation of Finite Difference equations in section 4.1. Specifically, the moving window implies a stencil of $S$ samples, $s_{0},s_{1},...,s_{S-1}]$, e.g. [1,0,1][-1,0,1] to sample 3 equispaced points centered on a point of interest. We then wish to find a $d^{\text{th}}$ order polynomial of the form $c_{0}+c_{1}s+c_{2}s^{2}+...+c_{d}s^{d}$ that when sampled at the stencil-points approximates the data, $\mathbf{y}_{\text{window}}$, at the corresponding points. Paralleling Equations 2 and 39, this can be set up as:
@@
$\frac{d}{dt}(c_{0}+c_{1}s+...+c_{d}s^{d})=\frac{d}{ds}(c_{0}+c_{1}s+...+c_{d}s^{d})\frac{ds}{dt}=\frac{c_{1}}{\Delta t}$
@@ 7.4.3 Spline Smoothing
where $S$ is the overall spline, $\boldsymbol{\theta}$ represents all coefficients collected together, and $\lambda$ is a hyperparameter which balances fitting the data with keeping the second derivative small. The second possibility [dierckx1975] is to simply require the sum of squares of the best fit be within a bound, $s$, which corresponds to the minimization of Equation 40 with $\lambda=0$ and an additional condition:
@@ 7.5.2 Radial Basis Functions
Local functions with independent variable given by a radial distance, $r$, which evaluate to zero beyond some maximum distance, $\rho$—such as truncated Gaussians, $e^{-r^{2}/(2\sigma^{2})},\ r<\rho$—can also constitute a linearly independent basis, making the matrix in the linear inverse problem of Equation 5 banded, symmetric, and invertible in $O(N\rho^{2})$. We call this the $\mathbf{A}$ matrix, and visualize an example in the first panel of Figure 42. Unfortunately, $\mathbf{A}$ is typically ill-conditioned, having eigenvalues very close to 0, which makes coefficient solutions highly sensitive to noise in the target, $\mathbf{y}$. Unlike the Fourier case where sensitive higher modes, corresponding to tiny eigenvalues of $\mathbf{A}$, can be zeroed out, discarding any element of a local basis significantly deforms function reconstruction in some region, so we need to keep them all.
@@
where $\lambda$ is a chosen parameter. By taking the gradient w.r.t. $\mathbf{c}$ and setting equal to 0, the above is minimized by solving the linear inverse problem $(\mathbf{A}^{T}\mathbf{A}+\lambda\mathbb{I})\mathbf{c}=\mathbf{A}^{T}\mathbf{y}$. We can eigendecompose symmetric $\mathbf{A}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{T}$ with $\mathbf{V}^{T}\mathbf{V}=\mathbb{I}$, and manipulate the solution into $(\boldsymbol{\Lambda}+\boldsymbol{\Lambda}^{-1}\lambda)\mathbf{c}^{\prime}=\mathbf{y}^{\prime}$, where $\mathbf{c}^{\prime}=\mathbf{V}^{T}\mathbf{c}$ and $\mathbf{y}^{\prime}=\mathbf{V}^{T}\mathbf{y}$ are just vector rotations. Because the matrices are diagonal, we can find $c_{i}^{\prime}=y_{i}^{\prime}/(\lambda_{i}+\lambda/\lambda_{i})$, where $\lambda_{i}$ is the $i^{\text{th}}$ eigenvalue from $\boldsymbol{\Lambda}$. In words, this means that for large eigenvalues, rotated coefficients are approximately proportional to rotated data divided by those eigenvalues, and as eigenvalues tend toward zero, coefficients do likewise.
@@
To see what this is doing relative to Tikhonov, we again eigendecompose to find $(\boldsymbol{\Lambda}+\mathbb{I}\lambda)\mathbf{c}^{\prime}=\mathbf{y}^{\prime}$, which has component-wise solution $c_{i}^{\prime}=y_{i}^{\prime}/(\lambda_{i}+\lambda)$. In words, large eigenvalues cause coefficients to be inversely proportional as before, but as $\lambda_{i}\to 0$, $c_{i}^{\prime}\to\frac{1}{\lambda}$ rather than 0 as for Tikhonov or $\infty$ in the undamped case.
@@
$\hat{x}(t)=\sum_{k=0}^{N-1}c_{k}\xi_{k}(t),\quad\xi_{k}(t)=\begin{cases}e^{-\frac{(t_{k}-t)^{2}}{2\sigma^{2}}}&|t_{k}-t|<\rho\\ 0&\text{otherwise}\end{cases}$
@@ 7.8.2 Accuracy and Bias
The other major consideration is method accuracy in terms of RMSE (Equation 33) and bias in terms of Error Correlation (Equation 34) against the true, withheld derivative. We are interested in whether the choice of method should change with different data characteristics: outliers, noise type, noise level, and the size of (consistent) $\Delta t$. We are also interested in whether the idiosyncrasies of particular data series favor particular differentiation algorithms, and we would like to know how sensitive methods are to the choice of cutoff frequency (bandlimit), which affects the smoothness of the optimized answers through the loss function’s $\gamma$, set according to the heuristic in Equation 37. In the presence of outliers, we optimize the robust loss function in Equation 36 with Huber parameter $M=2$, and in the absence of outliers, choose M=6M=6.282828We have observed empirically that, under conditions of Gaussian noise, optimizing the robust loss with M=6M\!=\!6 essentially always produces identical answers to optimizing the original RMSE-based loss, Equation 35. This makes sense theoretically, because inlier errors beyond $6\sigma$ from the mean error are exceedingly rare: 2*(1-scipy.stats.norm.cdf(6)) = 1.97e-9. Lowering $M$ to 2 provides a slight performance edge in the presence of outliers, allowing them to be weighted lighter, but in the absence of outliers M=2M\!=\!2 does begin to degrade performance—yet still only very slightly, because the portion of inliers within $2\sigma$ is more than 95%.
@@
The analysis space now has 7 dimensions: differentiation method, simulation, step size, noise type, noise scale, whether there are outliers, and cutoff frequency. There is also a bonus dimension: To get a sense of performance center and scatter, we need several rounds of simulation followed by optimization with different random seeds. Some dimensions must be explored exhaustively: 12 methods, 6 simulations, and we turn the number of random seeds all the way up to 52. To save compute and target our analysis at specific questions, we sweep along slices of the remaining orthotope, only varying one parameter at a time. We choose slices to all pass through a common central point, {cutoff frequency=3 Hz, $\Delta t$=0.01, noise type=normal, noise scale=1, outliers=False}), as symbolized in Figure 51.
@@ 7.9 Recommendations
To further compare methods, auxiliary considerations are summarized in Figure 57. These factors emphasize the value of algorithm flexibility, which can be underappreciated in theoretical discussion but is often important in practice. Indeed, one of the reasons neural networks have become so dominant is their endless capacity for recombination into architectures tailored to multivariate, multidimensional, sequential, incomplete, and structured data.303030“Time cannot wither nor custom stale her infinite variety.” –Shakespeare on the perceptron We recommend RTSDiff (section 7.7) as a general-purpose method for its outstanding versatility and narrowly superior accuracy in many experiments.
@@
Figure 57: Ability of methods to handle auxiliary considerations. Black tildes mean neither yes nor no, often that a method could be extended to handle a situation, but at the cost of computational efficiency or mathematical or code simplicity. The heavy dependency at issue is a set of convex solvers, which, although easy to install and use in modern form [cvxpy], are based on underlying compilers and iterative algorithms that can fail to solve efficiently or at all, a source of potential brittleness.
@@ 8 Irregularly-Spaced Samples
Sliding Polyfit $\checkmark$ Use sliding window in index space rather than dependent-variable space to capture same-sized groups of points across the domain.
@@ 8.1 Finite Difference with Irregular Steps
There is nothing stopping us from solving the stencil Vandermonde inverse problem of Equation 2 with irregular stencils. We do this by multiplying both sides by $\Delta x$ to put stencil locations back in units of the independent variable and then using whichever fractional or decimal stencil locations, $s_{0},...,s_{S-1}]$, match data locations.
@@ 8.3 Kalman Smoothing with Irregular Steps
$=\mathbb{E}\big[\big(e^{\mathcal{A}\Delta t}\mathbf{x}(0)+\cancel{\mathbf{B}(\Delta t)\mathbf{u}}+\mathbf{w}_{d}(\Delta t)-e^{\mathcal{A}\Delta t}\mathbb{E}[\mathbf{x}(0)]-\cancel{\mathbf{B}(\Delta t)\mathbf{u}}-\mathbb{E}[\mathbf{w}_{d}(\Delta t)]\big)\big(\text{ditto}\big)^{T}\big]$